# RoboArmGS: High-Quality Robotic Arm Splatting via Bézier Curve Refinement

### v1

@@ Abstract
Building high-quality digital assets of robotic arms is crucial yet challenging for the Real2Sim2Real pipeline. Current approaches naively bind static 3D Gaussians according to URDF links, forcing them to follow an URDF-rigged motion passively. However, real-world arm motion is noisy, and the idealized URDF-rigged motion cannot accurately model it, leading to severe rendering artifacts in 3D Gaussians. To address these challenges, we propose RoboArmGS, a novel hybrid representation that refines the URDF-rigged motion with learnable Bézier curves, enabling more accurate real-world motion modeling. To be more specific, we present a learnable Bézier Curve motion refiner that corrects per-joint residuals to address mismatches between real-world motion and URDF-rigged motion. RoboArmGS enables the learning of more accurate real-world motion while achieving a coherent binding of 3D Gaussians across arm parts. To support future research, we contribute a carefully collected dataset named RoboArm4D, which comprises several widely used robotic arms for evaluating the quality of building high-quality digital assets. We evaluate our approach on RoboArm4D, and RoboArmGS achieves state-of-the-art performance in real-world motion modeling and rendering quality. The code and dataset will be released.
@@ 1 Introduction
Robotic arms play a crucial role as the primary executors of tasks in contemporary automation and intelligent systems, making them a significant area of focus in the fields of computer vision and robotics [liang2024rapid, ma2024hierarchical, sun2025robava, jiang2024robots]. Reconstructing high-fidelity, interactable digital assets of robotic arms from real-world observations is a foundational step for advancing the Real2Sim2Real pipeline, which is critical for applications in robot control, policy learning, and system monitoring [wang2025embodiedreamer, pfaff2025scalable, abou2025real]. By generating these digital twins from simple video inputs, we can build simulation environments that faithfully mirror their real-world counterparts. This capability is crucial for bridging the notorious sim-to-real gap, enabling the development and validation of robust control policies in a safe and scalable manner [wagenmaker2024overcoming, jiang2024transic]. Consequently, the challenge of automatically generating these dynamic, high-fidelity assets from video remains a key bottleneck in robotics [xie2025vid2sim, jiang2025phystwin]. Overcoming this barrier is crucial for unlocking the vast potential of simulation. It enables large-scale generation of photorealistic, kinematically accurate data, which is essential for training the next generation of general-purpose robotic policies.
Current approaches for building dynamic digital assets of robotic arms typically involve complex, multi-stage pipelines, which hinder their scalability and robustness [lou2025robo, han2025re, li2024robogsim]. This pipeline typically begins with a strictly controlled data capture phase, requiring multi-view images or videos of the scene. Subsequently, the robotic arm must be precisely segmented from the background and other objects, a step that relies on complex segmentation algorithms or, in some cases, even manual annotation of point clouds [yang2025novel, lou2025robo]. This multi-stage pipeline is not only constrained by impractical data capture requirements that limit its use outside the lab but is also highly brittle; slight errors at any stage, particularly in coordinate frame alignment or part segmentation, can cause catastrophic failures in the final model’s fidelity. In contrast, learning directly from a single, casually captured monocular video offers a far more practical and scalable paradigm [li2025scalable, tao2025robopearls]. While this approach drastically lowers the data collection barrier, it introduces the deeper challenge of accurately modeling the arm’s complex and often non-ideal motion from unstructured observations. This is a problem that rigid, pre-defined kinematic models like Universal Robot Description Format (URDF) struggle to address.
@@
To address these challenges, we propose RoboArmGS (Fig. LABEL:fig:teaser), which introduces a novel hybrid representation that refines the URDF-rigged motion with learnable Bézier curves, enabling more accurate real-world motion modeling. Central to our approach are two key insights: (1) Instead of treating the 3D Gaussians as an unstructured whole, a structured binding mechanism that links them to a geometric prior, such as a mesh, can fundamentally enforce topological consistency. (2) The residual between an idealized kinematic model like URDF and real-world dynamics can be explicitly modeled and corrected using a flexible, learnable function. Our architecture is built upon these two principles, featuring a novel Gaussian binding strategy and a motion refiner.
Specifically, we first introduce our Structured Gaussian Binding (SGB), which binds each Gaussian to a specific face on the robot’s mesh. To be more precise, each Gaussian is parameterized within the local coordinate frame of its corresponding face, allowing for the learning of its color, opacity, and fine-grained positional offsets. This design ensures that the Gaussians inherit the underlying geometric structure of the robot, while also maintaining the flexibility to capture intricate, real-world appearances. For motion modeling, we then introduce our Bézier-based Motion Refiner (BMR). This module implements a hybrid approach where the URDF-rigged motion provides an idealized motion for the attached Gaussians. On top of this, the BMR learns a continuous-time residual transformation for each joint to precisely compensate for the residuals between the idealized motion and real-world motion. By integrating these two components, RoboArmGS allows for the generation of a statically consistent model while also accurately driving its dynamic motion. This approach significantly reduces rendering artifacts and results in a digital asset with high fidelity in both appearance and motion.
To validate our approach and to provide a standardized benchmark for future research in this area, we further contribute a new, carefully collected dataset, dubbed RoboArm4D. This dataset features several widely used robotic arms and is designed to comprehensively evaluate the quality of high-fidelity digital asset creation.
  • We propose RoboArmGS, a novel hybrid representation that uniquely combines Structured Gaussian Binding (SGB) for geometric consistency with a Bézier-based Motion Refiner (BMR) to accurately model real-world motions and mitigate rendering artifacts.
  • We introduce RoboArm4D, a new public benchmark dataset specifically designed for the task of creating high-fidelity digital assets of robotic arms from video.
  • We conduct extensive experiments demonstrating that RoboArmGS achieves state-of-the-art performance on the RoboArm4D benchmark in both rendering quality and motion accuracy.
@@ 2 Related Work
3D Gaussian Splatting (3DGS) has become a widely adopted representation for high-fidelity scene modeling [wei2025emd, wei2025gazeGaussian, wang2025plgs]. To capture non-rigid dynamics, Deformable 3DGS [yang2024deformable] introduces a deformation field to model monocular temporal variations. Zhang et al. [zhang2024dynamic, wei2025graphavatar] organize subsets of Gaussian points into graph structures and train graph neural networks to learn object dynamics under physical interactions. In avatar modeling, GaussianAvatars [qian2024gaussianavatars, chen2025mixedgaussianavatar] achieve precise parametric control of highly realistic facial avatars by binding discrete 3D Gaussians to a continuous FLAME mesh, jointly optimizing their parameters. Animatable Gaussians [li2024animatable] extend this Gaussian–mesh binding to full-body avatars using the SMPL template. In the domain of urban scene reconstruction, S3Gaussian [huang2024s3Gaussian] and StreetGaussians [yan2024street] achieve high-quality rendering and scene editing by separating Gaussians of moving cars and static background, while HUGS [zhou2024hugs] extends Gaussian splatting to additional modalities such as optical flow for improved dynamic modeling. 4DGF [fischer2024dynamic] adopts a hybrid neural scene representation, combining 3D gaussians as geometric supports with neural fields. OmniRe [chen2024omnire] constructs a dynamic neural scene graph based on 3DGS, generating multiple Gaussian representations for distinct dynamic participants and models pedestrians with the SMPL model. BézierGS [ma2025b] introduces learnable Bézier curves to represent vehicle motion trajectories, enforcing geometric consistency via an inter-group curve loss and reducing dependency on precise annotations. Compared to autonomous reconstruction methods that use Bézier curves for naive car motion representation, we employ a Bézier-curve-based correction module to refine an idealized forward kinematics model, aligning it more closely with real-world dynamics.
@@
3D reconstruction serves as a fundamental component for robotic policy learning and data synthesis [lu2024manigaussian, yu2025manigaussian++, chai2025gaf, wang2025embodiedocc++]. Recent research has increasingly leveraged 3DGS to enhance visual realism and geometry-aware robotic simulation. SplatSim [qureshi2025splatsim] replaces traditional mesh-based rendering with 3DGS within simulators, enabling photorealistic RGB data generation for manipulation policy training. RoboGSim [li2024robogsim] utilizes 3DGS to reconstruct scenes and objects from multi-view videos, creating mesh-based digital twins for simulation in Isaac Sim, facilitating demonstration synthesis and closed-loop evaluation. RE³SIM [han2025re] combines 3DGS for background rendering with mesh-based foreground modeling from multi-view stereo while RoboSplat [yang2025novel] employs a unified 3D Gaussian representation to model the entire workspace (robotic arm, target objects, and background) from multi-view images. More recently, Real2Render2Real [yu2025real2render2real] scans object with phones, converts 6-DoF object motion trajectories from human demonstration videos into executable robot actions, and leverages IsaacLab for scalable, photorealistic rendering, eliminating the need for simulation and physical hardware. ManipDreamer3D [li2025manipdreamer3d, li2025manipdreamer] represents scenes via occupancy from monocular images and synthesizes high-fidelity videos conditioned on optimized 3D trajectories. Robo-GS [lou2025robo] introduces a Gaussian–Mesh–Pixel binding technique to reconstruct static robotic arms and generate operable URDF assets from monocular videos, albeit with strong reliance on accurate panoramic annotations.
In contrast, we binding Gaussian primitives to robot arm meshes and focus on modeling realistic and precise robotic arm motions without requiring panoramic supervision.
@@ 3 Methodology
In this section, we introduce RoboArmGS (Fig. 2), our framework for creating high-quality digital twins of robotic arms. Our approach improves 3D Gaussian Splatting through two main contributions: Structured Gaussian Binding, which facilitates physically plausible articulated motion, and a Bézier-based Motion Refiner, designed to correct motion residuals. We also detail the optimization and regularization strategies employed to train our model.
Figure 2: Overview of RoboArmGS. Our method consists of two key modules: (a) Structured Gaussian Binding (SGB) binds 3D Gaussians according to the URDF links with adaptive densification; (b) Bézier-based Motion Refinement (BMR) corrects URDF-rigged motion and real-world motion discrepancies through learnable Bézier curves, enabling accurate novel-pose synthesis. The framework processes multi-view RGB images with known camera poses and robot joint angles, optimizing both modules jointly to achieve photorealistic rendering from novel viewpoints and robot motions.
@@ 3.1 Preliminaries
3D Gaussian Splatting [kerbl20233d] is a real-time radiance field method that represents a scene as a collection of 3D Gaussians. Each Gaussian is parameterized by a position $\mu$, rotation $q$, anisotropic scaling $s$, opacity $\alpha$, and Spherical Harmonics for color. During optimization, a valid positive semi-definite covariance matrix $\Sigma$ is constructed from the learnable scaling and rotation components. To render an image, these 3D Gaussians are projected onto the 2D image plane, sorted by depth, and then composited to compute the final pixel color $C$ via alpha blending:
@@
A Bézier curve [mortenson1999mathematics] is a parametric curve defined by a set of $n+1$ control points $\{P_{i}\}_{i=0}^{n}$. A point on the curve is computed as a weighted sum of these control points using Bernstein basis polynomials as weights:
@@
As $t$ varies from 0 to 1, the curve smoothly interpolates the control points, making it an effective tool for modeling smooth trajectories.
@@ 3.2 Structured Gaussian Binding
Unlike existing methods that naively bind static 3D Gaussians to URDF links, our Structured Gaussian Binding (SGB) module explicitly binds Gaussians to the kinematic mesh through physics-driven local parameterization and binding-aware adaptive densification. This mesh-level binding ensures motion consistency and structural coherence, effectively integrating kinematic constraints for high-fidelity motion representation.
@@
Inspired by recent work [qian2024gaussianavatars, xu2024gaussian], we bind each 3D Gaussian to a face on the robot’s kinematic mesh and drive its motion using a physics simulator. Our approach decomposes each Gaussian’s properties into dynamic local-to-world properties and static, learnable properties. The dynamic properties, which include its origin $T_{i}(t)$, orientation $R_{i}(t)$, and scaling factor $k_{i}(t)$, are derived from the mesh vertices after they are transformed by link poses from the MuJoCo engine [todorov2012mujoco]. The static properties consist of the learnable local attributes ($\mu_{j},r_{j},s_{j}$, color, and opacity), which are combined with the dynamic properties for rendering:
@@
$\mu^{\prime}_{j}(t)$ $=k_{i}(t)\cdot(R_{i}(t)\mu_{j})+T_{i}(t)$ (4)
$s^{\prime}_{j}(t)$ $=k_{i}(t)\cdot s_{j}$ (5)
This hybrid parameterization ensures physically-plausible motion via the rigid body structure, while simultaneously allowing for the capture of subtle geometric and appearance details through the optimization of local attributes.
@@
To represent high-frequency details, we adapt the density control from 3DGS [kerbl20233d] with two binding-aware modifications, similar to strategies for rigged avatars [qian2024gaussianavatars]. Binding inheritance ensures that new primitives created during densification inherit the parent face index from their progenitor. Additionally, a binding-preservation constraint modifies the standard pruning process to retain at least one Gaussian per face, which helps to prevent visual holes on occluded surfaces. These adaptations are essential for maintaining the model’s structural integrity during optimization.

### v2

@@ Abstract
Constructing photorealistic and controllable robotic arm digital assets from real observations is fundamental to robotic applications. Current approaches naively bind static 3D Gaussians according to URDF links, forcing them to follow an URDF-rigged motion passively. However, the idealized URDF-rigged motion cannot accurately model the actual motion captured in real-world observations, leading to severe rendering artifacts in 3D Gaussians. To address these challenges, we propose RoboArmGS, a novel hybrid representation that refines the URDF-rigged motion with learnable Bézier curves, enabling more accurate real-world motion modeling. To be more specific, we present a learnable Bézier Curve motion refiner that corrects per-joint residuals to address mismatches between real-world motion and URDF-rigged motion. RoboArmGS enables the learning of more accurate real-world motion while achieving a coherent binding of 3D Gaussians across arm parts. To support future research, we contribute a carefully collected dataset named RoboArm4D, which comprises several widely used robotic arms for evaluating the quality of building high-quality digital assets. We evaluate our approach on RoboArm4D, and RoboArmGS achieves state-of-the-art performance in real-world motion modeling and rendering quality. The code and dataset will be released.
Machine Learning, ICML
@@ 1 Introduction
Robotic arms play a crucial role as the primary executors of tasks in contemporary automation and intelligent systems, making them a significant area of focus in the fields of robotics (Liang et al., 2024; Ma et al., 2024; Sun et al., 2025; Jiang et al., 2024). Reconstructing high-fidelity, interactable digital assets of robotic arms from real-world observations is a foundational step for creating robust simulation platforms, which is critical for applications in robot control, policy learning, and system monitoring (Wang et al., 2025a; Pfaff et al., 2025; Abou-Chakra et al., 2025). By generating these digital assets from video inputs, we can build simulation environments that faithfully mirror their real-world counterparts. This capability is essential for bridging the sim-to-real gap, yet is fundamentally hindered by the discrepancy between idealized kinematic models and physical execution. While standard calibration techniques effectively rectify static geometric errors, they often fall short in capturing the complex dynamic deviations inherent in real-world motion. Consequently, the challenge of automatically generating dynamic, high-fidelity assets from video remains a key bottleneck in robotics (Xie et al., 2025; Jiang et al., 2025).
Current approaches for building dynamic digital assets of robotic arms typically involve complex, multi-stage pipelines, which hinder their scalability and robustness (Lou et al., 2025; Han et al., 2025; Li et al., 2024a). These pipelines usually begin with a strictly controlled data capture phase, requiring multi-view images or videos of the scene. Subsequently, the robotic arm requires precise segmentation to isolate it from the background and distinguish its individual articulated parts, a process that necessitates extensive labor-intensive manual annotation (Yang et al., 2025; Lou et al., 2025). Such requirements not only limit the method’s applicability outside laboratory settings but also introduce significant fragility. Crucially, this paradigm often assumes a perfect correspondence between the kinematic model and visual observations, thereby failing to account for dynamic discrepancies caused by real-world factors like control latency or joint friction.
In contrast, learning directly from a single, casually captured monocular video offers a more practical and scalable paradigm (Li et al., 2025a; Tao et al., 2025). While this approach drastically lowers the data collection barrier, it significantly intensifies the challenge of aligning the rigid kinematic prior with the actual, non-ideal motion observed in the footage. Addressing this requires accurately modeling these dynamic deviations to reconcile them with the standard Universal Robot Description Format (URDF), a step vital for enabling high-fidelity, motion-accurate simulation.
@@
To address these challenges, we propose RoboArmGS (Fig. LABEL:fig:_teaser), a novel hybrid representation designed to harmonize the rigid kinematic prior with actual visual observations. Our approach dynamically refines the idealized URDF-driven motion using a learnable Bézier curve model, enabling the creation of digital assets that are temporally coherent and precisely aligned with real-world footage. Central to our approach are two key insights: (1) Instead of treating 3D Gaussians as an unstructured collection, we introduce a structured binding mechanism that anchors them to a geometric prior, specifically the robot’s mesh, to fundamentally enforce topological consistency during motion. (2) The complex, time-varying discrepancies between the idealized URDF model and the physical robot’s dynamics can be explicitly captured and corrected using a flexible, smooth parameterization.
Guided by these insights, our architecture integrates two core modules. First, the Structured Gaussian Binding (SGB) anchors each Gaussian to a face on the robot’s mesh within its local coordinate frame. This ensures that the model inherits the robot’s rigid geometric structure while retaining the flexibility to capture intricate appearances via learnable offsets. Second, the Bézier-based Motion Refiner (BMR) addresses dynamic discrepancies by acting as a lightweight motion residual field. It leverages URDF kinematics as a strong motion prior and learns a smooth, continuous-time residual transformation to reconcile the rigid model with actual observations, thereby compensating for systematic errors in real-world motion. By synergizing these components, RoboArmGS produces digital assets that are both structurally consistent and kinematically accurate, significantly reducing rendering artifacts caused by motion mismatch.
To validate our approach and support future research, we present RoboArm4D, a specialized dataset featuring representative robotic arms. Designed as a rigorous benchmark for high-fidelity asset creation, it specifically targets the evaluation of how well the generated digital assets conform to real-world visual dynamics.
Our main contributions are as follows:
  • We propose RoboArmGS, which synergizes Structured Gaussian Binding (SGB) and Bézier-based Motion Refiner (BMR) to enforce geometric consistency while correcting dynamic discrepancies, thereby reconciling rigid kinematic priors with actual observations.
  • We contribute RoboArm4D, a benchmark tailored for robotic arm digitization. Featuring representative mainstream robotic arms with distinct morphologies, it evaluates rendering fidelity and motion accuracy against physical reality.
  • Extensive experiments on RoboArm4D demonstrate that RoboArmGS outperforms state-of-the-art methods, setting a new standard for photorealistic and physically aligned simulation.
@@ 2 Related Work
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has emerged as a dominant representation for high-fidelity scene modeling (Wei et al., 2025c, a; Wang et al., 2025d). To extend this to dynamic environments, Deformable 3DGS (Yang et al., 2024) introduces deformation fields to capture monocular temporal variations, while others (Zhang et al., 2024; Wei et al., 2025b) leverage graph neural networks to learn object dynamics under physical interactions. In the realm of articulated avatars, mesh-guided approaches have shown significant promise. For instance, GaussianAvatars (Qian et al., 2024; Chen et al., 2025) bind 3D Gaussians to the FLAME mesh for precise facial control, and Animatable Gaussians (Li et al., 2024b) apply similar binding strategies to SMPL-based bodies. These methods demonstrate the efficacy of anchoring Gaussians to a geometric prior, a principle we adapt for robotic structures. In urban scene reconstruction, methods like S3Gaussian (Huang et al., 2024b), StreetGaussians (Yan et al., 2024), and HUGS (Zhou et al., 2024) decompose dynamic entities (e.g., vehicles) from static backgrounds to achieve high-quality rendering and editing. More recently, BézierGS (Ma et al., 2025) utilizes learnable Bézier curves to parameterize the global motion trajectories of vehicles, enforcing geometric consistency via inter-group losses. While BézierGS employs curves to model the absolute motion path, our approach fundamentally differs by using Bézier curves to model the residual deviation from a kinematic prior. Specifically, we leverage the curves to refine an idealized forward kinematics chain, correcting the dynamic mismatch between the theoretical model and real-world observations.
@@
High-fidelity 3D reconstruction is pivotal for robotic policy learning and data synthesis (Lu et al., 2024; Yu et al., 2025b; Chai et al., 2025; Wang et al., 2025b). Recent approaches have integrated 3DGS to bridge the gap between simulation and reality. Several methods focus on integrating 3DGS into existing simulators: SplatSim (Qureshi et al., 2025) replaces mesh rendering with 3DGS for photorealistic output, while Real2Render2Real (Yu et al., 2025a) leverages scanned objects and human demonstrations to scale up training data generation in IsaacLab. However, for asset creation, most existing pipelines rely on complex multi-view setups. For instance, RoboGSim (Li et al., 2024a) and RE³SIM (Han et al., 2025) reconstruct scene assets from multi-view videos to build digital twins. Similarly, RoboSplat (Yang et al., 2025) uses a unified Gaussian representation from multi-view inputs to model the entire workspace. While effective, these multi-stage pipelines often suffer from scalability issues. On the other hand, monocular approaches like ManipDreamer3D (Li et al., 2025c, b) synthesize videos from single-view occupancy but focus less on precise kinematic control. Most relevant to our work is Robo-GS (Lou et al., 2025), which binds Gaussians to meshes to reconstruct robotic arms from monocular video. However, Robo-GS primarily targets static asset generation and relies on precise panoramic annotations. In contrast, our RoboArmGS not only utilizes monocular input without such heavy supervision but also explicitly models the dynamic discrepancies between the kinematic prior and real-world motion, ensuring both visual and kinematic fidelity during execution.
@@ 3 Methodology
In this section, we present RoboArmGS as shown in Fig. 2, a framework designed to bridge the sim-to-real motion discrepancy in high-fidelity robotic asset creation. Instead of relying solely on rigid kinematic assumptions, we introduce a hybrid representation that harmonizes explicit geometric priors with learnable dynamic refinements. Specifically, Structured Gaussian Binding (SGB) enforces topological consistency by anchoring Gaussians to the robot’s mesh, while the Bézier-based Motion Refiner (BMR) explicitly captures continuous residuals between idealized Foward Kinematic (FK) models and real-world observations. Finally, we detail the optimization and regularization strategies that ensure robust training from monocular video.
Figure 2: Overview of RoboArmGS. (a) SGB anchors Gaussians to URDF links with binding-aware densification for structural consistency. (b) BMR refines kinematic discrepancies via learnable Bézier curves. Jointly optimized from monocular input, our method enables photorealistic rendering for novel robot poses and viewpoints.
@@ 3.1 Preliminaries
3D Gaussian Splatting (Kerbl et al., 2023) is a real-time radiance field method that represents a scene as a collection of unstructured 3D Gaussians. Each Gaussian is parameterized by a position $\mu$, rotation $q$, anisotropic scaling $s$, opacity $\alpha$, and Spherical Harmonics for view-dependent color. During optimization, a valid positive semi-definite covariance matrix $\Sigma$ is constructed from the learnable scaling and rotation components. To render an image, these 3D Gaussians are projected onto the 2D image plane, sorted by depth, and then composited to compute the final pixel color $C$ via alpha blending:
@@
A Bézier curve (Mortenson, 1999) is a smooth, differentiable parametric curve defined by a set of $n+1$ control points $\{P_{i}\}_{i=0}^{n}$. A point on the curve is computed as a weighted sum of these control points using Bernstein basis polynomials as weights:
@@
where $B_{i,n}(t)=\binom{n}{i}t^{i}(1-t)^{n-i}$. As $t$ varies from 0 to 1, the curve provides a compact and continuous parameterization that smoothly interpolates within the convex hull of its control points. This makes it an effective tool for modeling temporally coherent trajectories and motion residuals.
@@ 3.2 Structured Gaussian Binding
Unlike existing methods that naively bind static 3D Gaussians to URDF links, our Structured Gaussian Binding (SGB) is designed to preserve the piecewise rigid structure of robotic arms. SGB explicitly binds each Gaussian to a specific face on the robot’s mesh, driven by physics-aware local parameterization and binding-aware adaptive densification. This mesh-level binding ensures all Gaussians on a single link undergo a collective rigid transformation, guaranteeing structural coherence and high-fidelity representation of the articulated motion.
@@
While inspired by recent advances in rigged avatars (Qian et al., 2024; Xu et al., 2024), we introduce a critical adaptation for rigid-body robotic systems. Unlike Linear Blend Skinning (LBS) (Li et al., 2017) models that often introduce non-rigid artifacts, we drive the motion using the robot’s forward kinematics to strictly maintain piecewise rigidity. Specifically, we bind each 3D Gaussian $j$ to a specific face $i$ on the kinematic mesh. This allows us to decompose the Gaussian properties into dynamic, pose-dependent states and learnable, static attributes. The dynamic properties consist of the local origin $T_{i}(t)$ and orientation $R_{i}(t)$, which are determined by the link poses obtained from the MuJoCo engine (Todorov et al., 2012) at time $t$. The static properties are defined as learnable local offsets that include position $\mu_{j}$, rotation $r_{j}$, scale $s_{j}$, color, and opacity. The final world-space state of a Gaussian for rendering is computed as:
$\mu^{\prime}_{j}(t)$ $=R_{i}(t)\mu_{j}+T_{i}(t)$ (3)
@@
$s^{\prime}_{j}(t)$ $=s_{j}$ (5)
where $\mu^{\prime}_{j}(t)$, $r^{\prime}_{j}(t)$, and $s^{\prime}_{j}(t)$ denote the world-space parameters. This ensures that all Gaussians anchored to the same link inherit a collective rigid transformation, while the learnable offsets $\mu_{j},r_{j}$ provide the degrees of freedom necessary to reconstruct complex surface geometries that exceed the resolution of the proxy URDF mesh.
@@
To capture high-frequency surface details while respecting the robot’s rigid structure, we adapt the density control mechanism of 3DGS (Kerbl et al., 2023) through two binding-aware modifications. First, we implement binding inheritance, ensuring that newly densified primitives automatically inherit the face index of their progenitor. This anchoring mechanism preserves piecewise rigid motion by preventing Gaussians from drifting away from their designated kinematic links during splitting or cloning. Second, we introduce a structural preservation constraint that prevents the pruning of the final Gaussian associated with any mesh face. This safeguard is critical for maintaining the geometric integrity of the robotic structure, effectively eliminating visual ”holes” that might otherwise emerge during large-scale articulated movements. Together, these adaptations ensure that the optimization process yields a high-fidelity representation that strictly adheres to the robot’s physical configuration and topological constraints.